OpenStudy (anonymous):

The probability of Andrew's team winning any given game in a 5-game series is 0.3. What is the probability that Andrew's team will win exactly 2 games in the series? 1) (.3)^2(.7)^3 2) 5(.3)^3(.7)^2 3) 10(.3)^2(.7)^3 4) 5(.3)^2(.7)

OpenStudy (anonymous):

as consistent to binomial distribution.

OpenStudy (anonymous):

sorry the answer is 3 I will show you why

OpenStudy (anonymous):

5 choose 2 which is 10 \[5!/2!(5-2)! \times 0.3^2\times0.7^3 = 10 \times 0.3^2 \times 0.7^3\]